in the experiment of caliberation of voltmeter a sandard cellof emf 1.1 volt is balanced against 440 cm of potentiometer wire. the potential difference across the ends of resistance is found to balance against 220 cm ofthe wire the corresponding reading of voltmeter is 0.5 vot.the error in the reading of voltmeter will be

Grade:12

2 Answers

Arun
4 years ago
If we have to find out the error of voltmeter if the same is showing 0.5V, then answer is D. Potential per unit length = 1.1/440=0.0025V/cm For length l=220 cm, voltmeter reading =0.0025x220=0.55V Error=0.5-0.55 =-0.05V
Vikas TU
4 years ago
Dear student 
E = 1.1 v 
L = 440 cm 
l= 220cm 
k = E/L 
k = 0.0025 Vm^-1
Vr = 0.0025 * 220 
Vr = 0.55
Voltmeter shows 0.5 v 
0.5-Vr = 0.5-0.55 = -0.05